OpenStudy (anonymous):

convert to vertex form y=x^2-2x-9 i got y=(x-1)^2 - 9, is this right?

OpenStudy (karon86):

oh thats easy

OpenStudy (catlover5925):

Vertex form is Y = a (x - h)^2 + k can you solve it now?

OpenStudy (karon86):

okay so translate \[y=ax ^{2}+ bx +c into y=a(x-h)^{2}+k\]

OpenStudy (karon86):

I mean into haha

OpenStudy (anonymous):

yeah i know what vertex form, i was jus kinda checking my answer @Catlover5925

OpenStudy (karon86):

y=ax^2+bx+c into y=a(x-h)+k a=1, b=-2, and c=-9 but you have to find the vertex. (h,k) or (x,y) in this case in order to find the the x part of the vertex, use the form. \[x=\left(\begin{matrix}-b /\ 2a\end{matrix}\right)\]

OpenStudy (karon86):

so substitute. -(-2)/2(1)

OpenStudy (anonymous):

So is my answer right? I do mines a different way by completing the square

OpenStudy (karon86):

So x=2/2 which =1 We just found the x part for the vertex (1,y) well. In order to find y you have to use that quadratic equation to get y So y=x^2-2x-9 right? use that: y=1(1)^2-2(1)-9 y=-10 you just found the y part of the vertex. The vertex is (1,-10) which is (h,k) plug (h,k) into the y=a(x-h)^2+k so y=a(x-1)^2-10 There, But there is one more thing you have to do. You have to plug in a=1 so y=1(x-1)-10. y=(x-1)^2-10 is the answer

OpenStudy (karon86):

Hope this helped. I am glad to help :)

OpenStudy (anonymous):

hm i was off by 1. can you double check your answer? i dont want to be wrong, but just want to be sure.. @wio

OpenStudy (anonymous):

nvm i understand now